OpenStudy (anonymous):

(3/4-2/3)+(1/2+1/3)=

OpenStudy (agreene):

(3/4 - 2/3) + (1/2 + 1/3) =(1/12) + (5/6) =11/12

OpenStudy (anonymous):

the same answer i got but this answer key says different thank you

OpenStudy (anonymous):

(3/4 - 2/3) + (1/2 + 1/3) =(1/12) + (5/6) =11/12

OpenStudy (anonymous):

yeah you're right